OpenStudy (anonymous):

Solve the following system of equations. 2x – 3y + z = –1 2x – 2y – 2z = –8 x + 5y – 3z = 2

OpenStudy (anonymous):

can you use matrix (linear algebra)? if not: 1) subtract second from first 2) multiply third eq by (-2) and add to first you'll get 2 equation with 2 variables - solve it need help with it?

OpenStudy (anonymous):

i thought u were supose to add the first and the second equations

OpenStudy (anonymous):

you can add, but in order to cancel out x you need to multiply one of the equations by (-1)

OpenStudy (anonymous):

you can re-write it as: 2x-3y+z=-1 -2x+2y+2z=8 -2x-10y+6z=-4 now add second to first - one equation add third to first - second. solve it for y & z

OpenStudy (anonymous):

I got: z=3, y=2, x=1

OpenStudy (anonymous):

i got that too. thanks tho

OpenStudy (anonymous):

it's nice the answer matching :) welcome!